OpenStudy (perl):

Hannah has $175 to spend. She buys $120 worth of non-taxable items. Some other items are taxable at 6%. Which inequality shows how much she can spend on taxable items before tax is applied? F. x<=$3.30 H. x<=$51.89 G. x<=$45.09 J. x<=$165.09

OpenStudy (mertsj):

120+1.06x<175

OpenStudy (perl):

can you explain

OpenStudy (whpalmer4):

non-taxable expenditure is $120 taxable is $1.06x ($1x for the base amount, plus $0.06x for the tax) she can spend up to $175, so the inequality is $120 + $1.06x <= $175 now just rearrange to match one of your answer choices

OpenStudy (perl):

I did 175 - 120 = 55 and then 55 - .06*55 = 51.7

OpenStudy (whpalmer4):

you have to do $55/1.06 = max you can spend to account for the tax correctly.

OpenStudy (whpalmer4):

it's x + 6%*x = $55 1.06x = $55 x = $55/1.06